OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 2, y = t2 - 4

OpenStudy (jdoe0001):

\(\bf x = t^2+2\qquad \qquad y = t^2-4\\ \quad \\ \quad \\ \textit{solving the 1st one for "t", we get}\\ x = t^2+2\implies \sqrt{x-2}= t\\ \quad \\ \textit{now that we know what "t" is, we substitute on the 2nd one}\\ y = t^2-4\implies y = \left(\sqrt{x-2}\right)^2-4\)

OpenStudy (jdoe0001):

and if you expand the right-side, you'd get your rectangular equation